# ,  Wei Jin

### v1

@@ 1 Introduction
We consider finite, simple, and undirected graphs and follow the terminology in [1]. The study of graph substructures and their enumeration has been a theme in extremal graph theory, dating back to a question posed by Erdős and Moser in 1960: what is the maximum possible number of maximal independent sets in general graphs of order $n$, and what are the graphs that achieve this maximum value? This was solved by Moon and Moser [6] in 1965. Since then, the maximum number of maximal or maximum independent sets on various graph families such as trees, unicyclic graphs, connected graphs and the graphs with at most $r$ cycles have been determined, refer [13, 3, 5, 7, 14, 18] for detail.
A subset $S$ of vertices in a graph $G$ is called a dissociation set if it induces a subgraph with vertex degree at most one. This concept, introduced by Yannakakis [15], not only generalizes the notions of independent sets and induced matchings but also forms a dual relationship with the 3-path vertex cover problem [2]. In recent years, scholars have begun to pay attention to determining the maximum (or minimum) number of maximal or maximum dissociation sets on various graph families and characterize the corroding extremal graphs such as trees [11, 10, 12], forests [8], unicyclic graphs [17], general and triangle-free graphs [9].
Recently, Yuan et al. [16] focused on the problem of the maximum number of all dissociation sets and the characterization of extremal graphs. They established the upper bound of the maximum number of dissociation sets among all connected graphs of order $n$ and characterized the corresponding extremal graphs. They also proposed a question regarding the second-largest number of dissociation sets and the corresponding extremal graphs as follows.
@@
$\begin{cases}K_{3}*\frac{n-3}{2}K_{2}\ or\ K_{1}*(2K_{1}\cup\frac{n-3}{2}K_{2})&\text{ if $n$ is odd; }\\ K_{3}*(\frac{n-4}{2}K_{2}\cup K_{1})\ or\ K_{1}*(3K_{1}\cup\frac{n-4}{2}K_{2})&\text{ if $n$ is even.}\end{cases}$
@@
In this paper, we give a positive answer to this question. The remainder of the paper is organized as follows: Section 2 introduces key definitions and preliminary lemmas; Section 3 presents the main results, including the upper bound for the second-largest number of dissociation sets and the corresponding extremal graphs.
@@ 2 Preliminaries
Let $G_{1},G_{2},\dots,G_{t}$ be all components of $G$. We can easily see that $d(G)=\prod_{i=1}^{i=t}d(G_{i})$. Two different vertices $u,v\in V(G)$ are called true twins if $N\left[u\right]=N\left[v\right]$ which indicates that $uv\in E(G)$ and are called false twins if $N(u)=N(v)$. It is easy to see that if $u,v$ are true twins in graph $G$, then $u,v$ are false twins in the subgraph $G-uv$.
A vertex with degree one is called a pendant vertex and the neighbor of a pendant vertex with degree at least two is called a quasi-pendant vertex. The dimension of cycle space is denoted by $c(G)=|E(G)|-|V(G)|+1$ where $|E(G)|$ and $|V(G)|$ represent the number of edges and vertices in $G$, respectively. Consider the cycle $C_{n}$ as $v_{1}v_{2}\cdots v_{n}v_{1}$. Denote by $C_{n}(k_{1},\cdots,k_{n})$ the unicyclic graph obtained from $C_{n}$ by identifying $v_{i}\in V(C_{n})$ with a vertex of pendant tree $T_{k_{i}}$ for $1\leq i\leq n$, where $k_{1}+\cdots+k_{n}$ equals to the order of $C_{n}(k_{1}\cdots k_{n})$.
@@
( [16]). Let $G$ be a graph with $v\in V(G)$. Then $d(G)\leq d(G-uv)$, the equality holds if and only if $N_{G}[u]=N_{G}[v]$.
@@
$F_{n}:=\begin{cases}K_{1}\ast\frac{n-1}{2}K_{2}&\text{ if $n$ is odd; }\\ P_{6}\ or\ K_{2}\ast 2K_{2}&\text{ if $n$ = 6; }\\ K_{2}\ast\frac{n-2}{2}K_{2}&\text{ if $n$ is even and $n$ $\neq$ 6.}\end{cases}$
@@
$h(n):=\begin{cases}2^{n-1}+(n+9)\cdot 2^{\frac{n-7}{2}}&\text{if }n\text{ is odd};\\ 2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}&\text{if }n\text{ is even and }n\neq 6;\\ 42&\text{if }n=6.\end{cases}$
@@
Let $G$ be a graph with quasi-pendant vertex $u_{q}$. Suppose that $v_{1},v_{2},\cdots,v_{s}$ are pendant vertices that are adjacent to $u_{q}$ where $s\geq 2$. Then $d(G)<d(G+v_{s-1}v_{s}-u_{q}v_{s})$.
Proof. Note that $v_{s-1},v_{s}$ are true twins in $G+v_{s-1}v_{s}$. By Lemma 2.3, $d(G)=d(G+v_{s-1}v_{s})$. Note that $u_{q},v_{s}$ are not true twins in the graph $G+v_{s-1}v_{s}$. By Lemma 2.3, $d(G+v_{s-1}v_{s})<d(G+v_{s-1}v_{s}-u_{q}v_{s})$. Then $d(G)<d(G+v_{s-1}v_{s}-u_{q}v_{s})$. $\Box$
@@
Proof. In contrast, suppose that $s\geq 4$. Choose four pendant vertices $v_{1},v_{2},v_{3},v_{4}$ that are adjacent to $u_{q}$. Note that $v_{1},v_{2}$ are true twins in $T+v_{1}v_{2}$. By Lemma 2.3, $d(T)=d(T+v_{1}v_{2})$. Let ${T}^{\prime}=T+v_{1}v_{2}-u_{q}v_{2}$ (see Figure 2). Note that $u_{q},v_{2}$ are not true twins in the graph $T+v_{1}v_{2}$. By Lemma 2.3, $d(T+v_{1}v_{2})<d(T+v_{1}v_{2}-u_{q}v_{2})=d(T^{\prime})$. Similarly, we can find that $d({T}^{\prime})=d({T}^{\prime}+v_{3}v_{4})<d({T}^{\prime}+v_{3}v_{4}-u_{q}v_{4})$. Let $T^{\prime\prime}={T}^{\prime}+v_{3}v_{4}-u_{q}v_{4}$. Consequently, $d(T)<d({T}^{\prime})<d(T^{\prime\prime})$ and $|V(T)|=|V(T^{\prime})|=|V(T^{\prime\prime})|$. This contradicts the fact that $T$ is the tree of order $n$ with the second-largest number of dissociation sets. So, $s\leq 3$. $\Box$
Figure 2: $T$, ${T}^{\prime}$, ${T}^{\prime\prime}$.
@@ 3 The second-largest number of dissociation sets in connected graph
[…]
@@
$T_{n}:=\begin{cases}K_{1}\ast(2K_{1}\cup\frac{n-3}{2}K_{2}),&\text{ if $n$ is odd ;}\\ K_{1}\ast(3K_{1}\cup\frac{n-4}{2}K_{2}),&\text{ if $n$ is even ,}\end{cases}$
@@
On the contrary, suppose $3\leq{\rm deg}_{T}(v_{1})\leq 4$. By the choice of $P$, all components of $T-v_{1}$ are isolated vertices except the component that contains $v_{2}$. Let $u$ be a pendant vertex that is adjacent to $v_{1}$ where $u\not=v_{0}$. Let ${\widetilde{T}=T+v_{0}u-v_{1}u}$, see Figure 3. By Lemma 2.6, $d(T)<d(\widetilde{T})$. Furthermore, $\widetilde{T}\ncong F_{n}$ (by Theorem 2.4). It follows that $d(T)<d(\widetilde{T})<f(n)$. This is contradictory because $T$ is the tree with the second-largest number of dissociation sets. Thus, ${\rm deg}_{T}(v_{1})=2$.
@@
In fact, suppose that $C$ is not an isolated vertex. By the choice of $P$, the diameter of the induced subgraph $T\left[V(C)\cup\left\{v_{2}\right\}\right]$ is no more than two. Thus, $C\cong K_{1,t}$. By Corollary 2.7, $t\leq 3$. To complete the proof of the claim, we need to show $t=1$. In fact, suppose $t\geq 2$. Choose a path $uvw$ from $C$. Clearly, $v$ is the center of $C$ which is adjacent to $v_{2}$ in $T$. By Lemma 2.6 and Theorem 2.4, $d(T)<d(T+wu-uv)<f(n)$. This contradicts the fact that $T$ is the tree with the second-largest number of dissociation sets. It follows that $C\cong K_{1,1}$.
@@
In contrast, suppose that $l>4$. Then $33\leq m\leq n-$ 3. By Lemma 2.2,
@@
[…]
  • $d(T-v_{0})=f(n-1)$.
@@
    $=2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}-(2^{n-6}-n\cdot 2^{\frac{n-8}{2}})$
    $<h(n)=2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}.$
    The last inequality follows from the fact that $02^{n-6}-n\cdot 2^{\frac{n-8}{2}}>$ 0 when $n\geq 12$. This is contradictory.
@@
    The last inequality follows from the fact that $02^{n-6}-(n-5)\cdot 2^{\frac{n-9}{2}}>$ 0 when $n\geq 11$. This is contradictory. Here, we have completed the proof of Claim $3$.
@@
In fact, suppose $d\geq 2$. Since $G$ is connected, we can choose $d-1$ edges $u_{1}v_{1},u_{2}v_{2},\ldots,u_{d-1}v_{d-1}$ of $G$ such that the subgraph
@@
Hence, $d(G)=h(n)$ if and only if all of the above inequalities hold. By Theorem 2.5, $d({G}^{\prime})=h(n)$ is equivalent to ${G}^{\prime}\cong U_{n}$. By Lemma 2.3, $d(G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2})=d({G}^{\prime})$ is equivalent to $u_{d-1},v_{d-1}$ are true twins in $G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2}$ It is follows that $u_{d-1},v_{d-1}$ are false twins in ${G}^{\prime}$. But this contradicts the fact that there are no false twins in ${G}^{\prime}\cong U_{n}$, where $U_{n}$ is defined in Theorem 2.5. Then $d(G)<h(n)$. $\Box$
@@
$\begin{array}[]{llll}d(T_{2}^{9})<d(T_{1}^{9}),&d(T_{7}^{9})<d(T_{3}^{9}),&d(T_{8}^{9})<d(T_{5}^{9}),&d(T_{9}^{9})<d(T_{6}^{9}),\vspace{1.0ex}\\ d(T_{13}^{9})<d(T_{10}^{9}),&d(T_{19}^{9})<d(T_{1}^{9}),&d(T_{20}^{9})<d(T_{3}^{9}),&d(T_{21}^{9})<d(T_{5}^{9}),\vspace{1.0ex}\\ d(T_{23}^{9})<d(T_{4}^{9}),&d(T_{24}^{9})<d(T_{4}^{9}),&d(T_{25}^{9})<d(T_{6}^{9}),&d(T_{28}^{9})<d(T_{3}^{9}),\vspace{1.0ex}\\ d(T_{29}^{9})<d(T_{5}^{9}),&d(T_{30}^{9})<d(T_{22}^{9}),&d(T_{31}^{9})<d(T_{6}^{9}),&d(T_{32}^{9})<d(T_{10}^{9}),\vspace{1.0ex}\\ d(T_{33}^{9})<d(T_{22}^{9}),&d(T_{36}^{9})<d(T_{10}^{9}),&d(T_{38}^{9})<d(T_{27}^{9}),&d(T_{39}^{9})<d(T_{27}^{9}),\vspace{1.0ex}\\ d(T_{40}^{9})<d(T_{26}^{9}),&d(T_{41}^{9})<d(T_{27}^{9}),&d(T_{43}^{9})<d(T_{26}^{9}),&d(T_{44}^{9})<d(T_{4}^{9}),\vspace{1.0ex}\\ d(T_{45}^{9})<d(T_{6}^{9}),&d(T_{46}^{9})<d(T_{26}^{9}),&d(T_{47}^{9})<d(T_{6}^{9}).\end{array}$
@@
$\begin{array}[]{llll}d(T_{2}^{10})<d(T_{1}^{10}),&d(T_{9}^{10})<d(T_{3}^{10}),&d(T_{10}^{10})<d(T_{5}^{10}),&d(T_{11}^{10})<d(T_{7}^{10}),\vspace{1.0ex}\\ d(T_{16}^{10})<d(T_{12}^{10}),&d(T_{17}^{10})<d(T_{13}^{10}),&d(T_{18}^{10})<d(T_{14}^{10}),&d(T_{27}^{10})<d(T_{1}^{10}),\vspace{1.0ex}\\ d(T_{28}^{10})<d(T_{3}^{10}),&d(T_{29}^{10})<d(T_{5}^{10}),&d(T_{31}^{10})<d(T_{4}^{10}),&d(T_{32}^{10})<d(T_{6}^{10}),\vspace{1.0ex}\\ d(T_{33}^{10})<d(T_{7}^{10}),&d(T_{36}^{10})<d(T_{4}^{10}),&d(T_{37}^{10})<d(T_{7}^{10}),&d(T_{39}^{10})<d(T_{8}^{10}),\vspace{1.0ex}\\ d(T_{44}^{10})<d(T_{3}^{10}),&d(T_{45}^{10})<d(T_{30}^{10}),&d(T_{46}^{10})<d(T_{5}^{10}),&d(T_{47}^{10})<d(T_{34}^{10}),\vspace{1.0ex}\\ d(T_{48}^{10})<d(T_{7}^{10}),&d(T_{49}^{10})<d(T_{35}^{10}),&d(T_{50}^{10})<d(T_{35}^{10}),&d(T_{51}^{10})<d(T_{12}^{10}),\vspace{1.0ex}\\ d(T_{52}^{10})<d(T_{38}^{10}),&d(T_{53}^{10})<d(T_{7}^{10}),&d(T_{54}^{10})<d(T_{41}^{10}),&d(T_{55}^{10})<d(T_{42}^{10}),\vspace{1.0ex}\\ d(T_{56}^{10})<d(T_{13}^{10}),&d(T_{57}^{10})<d(T_{43}^{10}),&d(T_{58}^{10})<d(T_{41}^{10}),&d(T_{60}^{10})<d(T_{14}^{10}),\vspace{1.0ex}\\ d(T_{63}^{10})<d(T_{12}^{10}),&d(T_{65}^{10})<d(T_{13}^{10}),&d(T_{67}^{10})<d(T_{59}^{10}),&d(T_{68}^{10})<d(T_{14}^{10}),\vspace{1.0ex}\\ d(T_{73}^{10})<d(T_{30}^{10}),&d(T_{74}^{10})<d(T_{35}^{10}),&d(T_{75}^{10})<d(T_{59}^{10}),&d(T_{76}^{10})<d(T_{43}^{10}),\vspace{1.0ex}\\ d(T_{77}^{10})<d(T_{41}^{10}),&d(T_{80}^{10})<d(T_{59}^{10}),&d(T_{83}^{10})<d(T_{8}^{10}),&d(T_{84}^{10})<d(T_{41}^{10}),\vspace{1.0ex}\\ d(T_{85}^{10})<d(T_{40}^{10}),&d(T_{86}^{10})<d(T_{42}^{10}),&d(T_{88}^{10})<d(T_{7}^{10}),&d(T_{89}^{10})<d(T_{34}^{10}),\vspace{1.0ex}\\ d(T_{90}^{10})<d(T_{38}^{10}),&d(T_{91}^{10})<d(T_{4}^{10}),&d(T_{92}^{10})<d(T_{6}^{10}),&d(T_{93}^{10})<d(T_{13}^{10}),\vspace{1.0ex}\\ d(T_{94}^{10})<d(T_{41}^{10}),&d(T_{95}^{10})<d(T_{7}^{10}),&d(T_{96}^{10})<d(T_{41}^{10}),&d(T_{97}^{10})<d(T_{42}^{10}),\vspace{1.0ex}\\ d(T_{98}^{10})<d(T_{87}^{10}),&d(T_{99}^{10})<d(T_{34}^{10}),&d(T_{100}^{10})<d(T_{38}^{10}),&d(T_{101}^{10})<d(T_{13}^{10}),\vspace{1.0ex}\\ d(T_{102}^{10})<d(T_{87}^{10}),&d(T_{103}^{10})<d(T_{41}^{10}),&d(T_{104}^{10})<d(T_{8}^{10}),&d(T_{105}^{10})<d(T_{40}^{10}).\end{array}$

### v2

@@ 1 Introduction
We consider finite, simple, and undirected graphs and follow the terminology in [1]. The study of graph substructures and their enumeration has been a theme in extremal graph theory, dating back to a question posed by Erdős and Moser in 1960: what is the maximum possible number of maximal independent sets in general graphs of order $n$, and what are the graphs that achieve this maximum value? This was solved by Moon and Moser [6] in 1965. Since then, the maximum number of maximal or maximum independent sets on various graph families such as trees, unicyclic graphs, connected graphs and the graphs with at most $r$ cycles have been determined ; see [13, 3, 5, 7, 14, 18] for details.
A subset $S$ of vertices in a graph $G$ is called a dissociation set if it induces a subgraph with vertex degree at most one. This concept, introduced by Yannakakis [15], not only generalizes the notions of independent sets and induced matchings but also forms a dual relationship with the 3-path vertex cover problem [2]. In recent years, scholars have begun to pay attention to determining the maximum (or minimum) number of maximal or maximum dissociation sets on various graph families and characterize the corresponding extremal graphs such as trees [11, 10, 12], forests [8], unicyclic graphs [17], general and triangle-free graphs [9].
Recently, Yuan et al. [16] focused on the problem of the maximum number of all dissociation sets and the characterization of extremal graphs. They established the maximum number of dissociation sets among all connected graphs of order $n$ and characterized the corresponding extremal graphs. They also proposed a question of determining the second-largest number of dissociation sets and the corresponding extremal graphs as follows.
@@
[…]
Here, $h(n)$ is defined as follows
[…]
@@
In this paper, we give a positive answer to this question. The remainder of the paper is organized as follows: Section 2 introduces key definitions and preliminary lemmas; Section 3 presents the main results, including the second-largest number of dissociation sets in connected graphs of order $n$ and the corresponding extremal graphs.
@@ 2 Preliminaries
Let $G_{1},G_{2},\dots,G_{t}$ be all components of $G$. We can easily see that $d(G)=\prod_{i=1}^{i=t}d(G_{i})$. Two different vertices $u,v\in V(G)$ are called true twins if $N\left[u\right]=N\left[v\right]$ which indicates that $uv\in E(G)$ and are called false twins if $N(u)=N(v)$. It is easy to see that if $u$ and $v$ are true twins in graph $G$, then $u$ and $v$ are false twins in the subgraph GuvG-uv.
A vertex with degree one is called a pendant vertex and the neighbor of a pendant vertex with degree at least two is called a quasi-pendant vertex. The dimension of cycle space of a connected graph $G$ is denoted by c(G)=|E(G)||V(G)|+1c(G)=|E(G)|-|V(G)|+1 where $|E(G)|$ and $|V(G)|$ represent the number of edges and vertices in $G$, respectively. Consider the cycle $C_{n}$ as $v_{1}v_{2}\cdots v_{n}v_{1}$. Denote by $C_{n}(k_{1},\cdots,k_{n})$ the unicyclic graph obtained from $C_{n}$ by identifying $v_{i}\in V(C_{n})$ with a vertex of pendant tree $T_{k_{i}}$ for $1\leq i\leq n$, where $k_{1}+\cdots+k_{n}$ equals to the order of $C_{n}(k_{1},\cdots,k_{n})$.
@@
( [16]). Let $G$ be a graph with $uv\in E(G)$. Then $d(G)\leq d(G-uv)$, the equality holds if and only if $N_{G}[u]=N_{G}[v]$.
@@
[…]
@@
[…]
@@
Let $G$ be a graph of order $n\geq 4$ with quasi-pendant vertex $u_{q}$. Suppose that $v_{1},v_{2},\cdots,v_{s}$ are pendant vertices that are adjacent to $u_{q}$ where $s\geq 2$. Then $d(G)<d(G+v_{s-1}v_{s}-u_{q}v_{s})$.
Proof. Note that $v_{s-1}$ and $v_{s}$ are true twins in $G+v_{s-1}v_{s}$. By Lemma 2.3, $d(G)=d(G+v_{s-1}v_{s})$. Note that $u_{q}$ and $v_{s}$ are not true twins in the graph $G+v_{s-1}v_{s}$. By Lemma 2.3, $d(G+v_{s-1}v_{s})<d(G+v_{s-1}v_{s}-u_{q}v_{s})$. Then $d(G)<d(G+v_{s-1}v_{s}-u_{q}v_{s})$. $\Box$
@@
Proof. In contrast, suppose that $s\geq 4$. Choose four pendant vertices $v_{1},v_{2},v_{3},v_{4}$ that are adjacent to $u_{q}$. By Lemma 2.6, $d(T)<d(T+v_{1}v_{2}-u_{q}v_{2}):=d(T^{\prime})$. Similarly, $d({T}^{\prime})<d({T}^{\prime}+v_{3}v_{4}-u_{q}v_{4}):=d(T^{\prime\prime})$. Consequently, $d(T)<d({T}^{\prime})<d(T^{\prime\prime})$ and $|V(T)|=|V(T^{\prime})|=|V(T^{\prime\prime})|$. This contradicts the fact that $T$ is the tree of order $n$ with the second-largest number of dissociation sets. So, $s\leq 3$. $\Box$
Figure 2: $T$, $^{\prime}, T′′{T}^{\prime\prime}$.
@@ 3 The second-largest number of dissociation sets in connected graph
Let $T$ be a tree of order n(n=9or 10)n(n=9\ or\ 10) where $T\notin\{K_{1}\ast 4K_{2},K_{2}\ast 4K_{2}\}$. Then $d(T)\leq h(|V(T)|)$, the equality holds if and only if $T\in\{K_{1}\ast(2K_{1}\cup 3K_{2}),K_{1}\ast(3K_{1}\cup 3K_{2})\}$ where $h(9)=292,h(10)=556$.
@@
[…]
@@
On the contrary, suppose $3\leq{\rm deg}_{T}(v_{1})\leq 4$. By the choice of $P$, all components of $T-v_{1}$ are isolated vertices except the component that contains $v_{2}$. Let $u$ be a pendant vertex that is adjacent to $v_{1}$ where $u\not=v_{0}$. Let $\widetilde{T}=T+v_{0}u-v_{1}u$}, see Figure 3. By Lemma 2.6, d(T)<d(T~)d(T)<d($\widetilde${T}). Furthermore, $\widetilde{T}\ncong F_{n}$ (by Theorem 2.4). It follows that $d(T)<d(\widetilde{T})<f(n)$. This is contradictory because $T$ is the tree with the second-largest number of dissociation sets. Thus, ${\rm deg}_{T}(v_{1})=2$.
@@
In fact, suppose that $C$ is not an isolated vertex. By the choice of $P$, the diameter of the induced subgraph $T\left[V(C)\cup\left\{v_{2}\right\}\right]$ is no more than two. Thus, $C\cong K_{1,t}$. By Corollary 2.7, $t\leq 3$. To complete the proof of the claim, we need to show $t=1$. In fact, suppose $t\geq 2$. Choose a path uvwuvw from $C$. Clearly, $v$ is the center of $C$ which is adjacent to $v_{2}$ in $T$. By Lemma 2.6 and Theorem 2.4, d(T)<d(T+wuuv)<f(n)d(T)<d(T+wu-uv)<f(n). This contradicts the fact that $T$ is the tree with the second-largest number of dissociation sets. It follows that $C\cong K_{1,1}$.
@@
In contrast, suppose that $l>4$. Then $3\leq m\leq n-3$. By Lemma 2.2,
@@
$\begin{aligned} d(T-v_{0}-v_{1}-v_{2})&\leq 2^{n-3-m}\cdot f(m)\leq 2^{n-6}\cdot f(3)=7\cdot 2^{n-6}\\ \end{aligned}.$
We have found that there are two different cases of $d(T-v_{0})$. Let us discuss them separately below.
Case 1: $d(T-v_{0})=f(n-1)$.
@@
[…]
The last inequality follows from the fact that $2^{n-6}-n\cdot 2^{\frac{n-8}{2}}>0$ when $n\geq 12$. This is contradictory.
@@
The last inequality follows from the fact that $2^{n-6}-(n-5)\cdot 2^{\frac{n-9}{2}}>0$ when $n\geq 11$. This is contradictory. Here, we have completed the proof of Claim $3$.
@@
In fact, suppose $d\geq 2$. Since $G$ is connected, we can choose d1d-1 edges $u_{1}v_{1},u_{2}v_{2},\ldots,u_{d-1}v_{d-1}$ of $G$ such that the subgraph
@@
Hence, $d(G)=h(n)$ if and only if all of the above inequalities hold. By Theorem 2.5, $d({G}^{\prime})=h(n)$ if and only if ${G}^{\prime}\cong U_{n}$. By Lemma 2.3, $d(G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2})=d({G}^{\prime})$ is equivalent to $u_{d-1}$ and $v_{d-1}$ are true twins in $G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2}$. It follows that $u_{d-1}$ and $v_{d-1}$ are false twins in ${G}^{\prime}$. But this contradicts the fact that there are no false twins in ${G}^{\prime}\cong U_{n}$, where $U_{n}$ is defined in Theorem 2.5. Then $d(G)<h(n)$. $\Box$
@@
$\begin{array}[]{llll}d(T_{2}^{9})<d(T_{1}^{9}),&d(T_{7}^{9})<d(T_{3}^{9}),&d(T_{8}^{9})<d(T_{5}^{9}),&d(T_{9}^{9})<d(T_{6}^{9}),\vskip 4.30554pt\\ d(T_{13}^{9})<d(T_{10}^{9}),&d(T_{19}^{9})<d(T_{1}^{9}),&d(T_{20}^{9})<d(T_{3}^{9}),&d(T_{21}^{9})<d(T_{5}^{9}),\vskip 4.30554pt\\ d(T_{23}^{9})<d(T_{4}^{9}),&d(T_{24}^{9})<d(T_{4}^{9}),&d(T_{25}^{9})<d(T_{6}^{9}),&d(T_{28}^{9})<d(T_{3}^{9}),\vskip 4.30554pt\\ d(T_{29}^{9})<d(T_{5}^{9}),&d(T_{30}^{9})<d(T_{22}^{9}),&d(T_{31}^{9})<d(T_{6}^{9}),&d(T_{32}^{9})<d(T_{10}^{9}),\vskip 4.30554pt\\ d(T_{33}^{9})<d(T_{22}^{9}),&d(T_{36}^{9})<d(T_{10}^{9}),&d(T_{38}^{9})<d(T_{27}^{9}),&d(T_{39}^{9})<d(T_{27}^{9}),\vskip 4.30554pt\\ d(T_{40}^{9})<d(T_{26}^{9}),&d(T_{41}^{9})<d(T_{27}^{9}),&d(T_{43}^{9})<d(T_{26}^{9}),&d(T_{44}^{9})<d(T_{4}^{9}),\vskip 4.30554pt\\ d(T_{45}^{9})<d(T_{6}^{9}),&d(T_{46}^{9})<d(T_{26}^{9}),&d(T_{47}^{9})<d(T_{6}^{9}).\end{array}$
@@
$\begin{array}[]{llll}d(T_{2}^{10})<d(T_{1}^{10}),&d(T_{9}^{10})<d(T_{3}^{10}),&d(T_{10}^{10})<d(T_{5}^{10}),&d(T_{11}^{10})<d(T_{7}^{10}),\vskip 4.30554pt\\ d(T_{16}^{10})<d(T_{12}^{10}),&d(T_{17}^{10})<d(T_{13}^{10}),&d(T_{18}^{10})<d(T_{14}^{10}),&d(T_{27}^{10})<d(T_{1}^{10}),\vskip 4.30554pt\\ d(T_{28}^{10})<d(T_{3}^{10}),&d(T_{29}^{10})<d(T_{5}^{10}),&d(T_{31}^{10})<d(T_{4}^{10}),&d(T_{32}^{10})<d(T_{6}^{10}),\vskip 4.30554pt\\ d(T_{33}^{10})<d(T_{7}^{10}),&d(T_{36}^{10})<d(T_{4}^{10}),&d(T_{37}^{10})<d(T_{7}^{10}),&d(T_{39}^{10})<d(T_{8}^{10}),\vskip 4.30554pt\\ d(T_{44}^{10})<d(T_{3}^{10}),&d(T_{45}^{10})<d(T_{30}^{10}),&d(T_{46}^{10})<d(T_{5}^{10}),&d(T_{47}^{10})<d(T_{34}^{10}),\vskip 4.30554pt\\ d(T_{48}^{10})<d(T_{7}^{10}),&d(T_{49}^{10})<d(T_{35}^{10}),&d(T_{50}^{10})<d(T_{35}^{10}),&d(T_{51}^{10})<d(T_{12}^{10}),\vskip 4.30554pt\\ d(T_{52}^{10})<d(T_{38}^{10}),&d(T_{53}^{10})<d(T_{7}^{10}),&d(T_{54}^{10})<d(T_{41}^{10}),&d(T_{55}^{10})<d(T_{42}^{10}),\vskip 4.30554pt\\ d(T_{56}^{10})<d(T_{13}^{10}),&d(T_{57}^{10})<d(T_{43}^{10}),&d(T_{58}^{10})<d(T_{41}^{10}),&d(T_{60}^{10})<d(T_{14}^{10}),\vskip 4.30554pt\\ d(T_{63}^{10})<d(T_{12}^{10}),&d(T_{65}^{10})<d(T_{13}^{10}),&d(T_{67}^{10})<d(T_{59}^{10}),&d(T_{68}^{10})<d(T_{14}^{10}),\vskip 4.30554pt\\ d(T_{73}^{10})<d(T_{30}^{10}),&d(T_{74}^{10})<d(T_{35}^{10}),&d(T_{75}^{10})<d(T_{59}^{10}),&d(T_{76}^{10})<d(T_{43}^{10}),\vskip 4.30554pt\\ d(T_{77}^{10})<d(T_{41}^{10}),&d(T_{80}^{10})<d(T_{59}^{10}),&d(T_{83}^{10})<d(T_{8}^{10}),&d(T_{84}^{10})<d(T_{41}^{10}),\vskip 4.30554pt\\ d(T_{85}^{10})<d(T_{40}^{10}),&d(T_{86}^{10})<d(T_{42}^{10}),&d(T_{88}^{10})<d(T_{7}^{10}),&d(T_{89}^{10})<d(T_{34}^{10}),\vskip 4.30554pt\\ d(T_{90}^{10})<d(T_{38}^{10}),&d(T_{91}^{10})<d(T_{4}^{10}),&d(T_{92}^{10})<d(T_{6}^{10}),&d(T_{93}^{10})<d(T_{13}^{10}),\vskip 4.30554pt\\ d(T_{94}^{10})<d(T_{41}^{10}),&d(T_{95}^{10})<d(T_{7}^{10}),&d(T_{96}^{10})<d(T_{41}^{10}),&d(T_{97}^{10})<d(T_{42}^{10}),\vskip 4.30554pt\\ d(T_{98}^{10})<d(T_{87}^{10}),&d(T_{99}^{10})<d(T_{34}^{10}),&d(T_{100}^{10})<d(T_{38}^{10}),&d(T_{101}^{10})<d(T_{13}^{10}),\vskip 4.30554pt\\ d(T_{102}^{10})<d(T_{87}^{10}),&d(T_{103}^{10})<d(T_{41}^{10}),&d(T_{104}^{10})<d(T_{8}^{10}),&d(T_{105}^{10})<d(T_{40}^{10}).\end{array}$